Kate makes a donation to a charitable organisation on a monthly basis. Her monthly donation is directly proportional to the square of her monthly savings. If she saves $900 and $1200 in January and February respectively, her donation increases by $35 from January to February. Find the amount of money she donates to the charitable organisation in each of the two months.
Answer by jorel555(1290)

Let n=the monthly proportion. Then:
900(n)+35=1200(n)
300n=35
n=35/300=11.67%!!!!!!!!!!!!!!!!